A sample of SiH4 gas has a volume of 42.3 L at 100ºC. The gas is cooled to -140ºC at constant pressure. What is the final volume of nitrogen?
Answer #1

At constant pressure the volume of an ideal gas is directly proportional to the absolute temperature of the gas.

Initial Temperature = 100 oC = (273+100) K = 373 K.

Final Temperature = -140 oC = (273-140) K = 133 K.

Initial Volume is 42.3 L.

Final Volume is calculated as

    { 42.3 L * (133 K / 373 K)} = 15.1 L.